OpenStudy (anonymous):

Find the value of Q in the following system so that the solution to the system is: {(x,y) : x-3y = 4} x-3y=4 Qx-6y=8

OpenStudy (anonymous):

Q=2

OpenStudy (anonymous):

how did you get that answer? @ask1709

OpenStudy (anonymous):

ok i looked at it and i saw that since its a system, the equations are the same. so, if you look at the equations, u see that the second equation is just the same as the first on but it is expressed 2 times larger

OpenStudy (anonymous):

-3y*2=-6y

OpenStudy (anonymous):

4*2=8

OpenStudy (anonymous):

so the answer would be Q=8?

OpenStudy (anonymous):

no it would be 2

OpenStudy (anonymous):

since 1*2=?

OpenStudy (anonymous):

sorry i'm not a good explainer

OpenStudy (anonymous):

its okay! are you still in 7th grade? @ask1709

OpenStudy (anonymous):

no now i just got in 8 but I now do geometry(10th grade math)

OpenStudy (anonymous):

oh wow you're good! Im doing pre-cal :p

OpenStudy (anonymous):

thanks @ali1029 but how come you are asking algebra 1 questions?

OpenStudy (anonymous):

youre welcome. & lol im in pre-cal but it just started and i guess we're reviewing algebra 1 &2 but i suck at math so i dont know anything lol @ask1709

OpenStudy (anonymous):

So you are 11th grade?

OpenStudy (anonymous):

Well I am in geometry but I can do basic pre cal. thats why I enjoy this website because i like solving math.

OpenStudy (anonymous):

thats awesome for you! i wish math was my good subject :( im better in english and all that. lol emberassing but im a senior this year @ask1709

OpenStudy (anonymous):

thanks @ali1029 but I tutor a junior and he is just like you he is really good at literature, but is confused about math

OpenStudy (anonymous):

yes that is definitely me all the way haha, well congrats to you! you're amazing for your age!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

But you are definitely a better explainer than me!

OpenStudy (anonymous):

haha no worries! i can only explain the super easy stuff, anything more advanced- forget it! :p @ask1709